OpenStudy (anonymous):

how to convert 45.62 kilojoules in to kilocalories ( 1 cal = 4.184 Joule)

OpenStudy (anonymous):

1cal=4.184j 1J=1/4.184 cal 45620 J = 45620/4.184cal=45.62/4.184 kcal

OpenStudy (king):

10.9 kcal